OpenStudy (anonymous):

Find the components of a vector with the initial point P_1(2,4) and the terminal point P_0(2,-9). a.) <4,13> b.) <-4,-13> c.) <0,-13> d.) <0,13>

OpenStudy (nurali):

Components of the vector = Terminal point - Initial point = 2,-9 - 2,4 = 0, -13 = Answer

OpenStudy (anonymous):

What do u kno so far be for we get started

OpenStudy (anonymous):

I thought this was a helping site